# Beyond LLMs, Sparse Distributed Memory, and Neuromorphics ¡A Hyper-Dimensional SRAM-CAM ”VaCoAl” for Ultra-High Speed, Ultra-Low Power, and Low Cost¿

### 13 Apr 2026 v1

@@ 1.1 Resolving the “Singularities” of Modern Generative AI and Transcending Kanerva’s Theory
Before turning to the structural limitations of modern AI that motivate this work, we state at the outset the central empirical finding that the rest of the paper is organised to explain. In Section 5 we report that a deterministic hyperdimensional computing architecture built on Galois-field algebra exhibits, on a real-world ontology of approximately 470,000 mentor–student records traversed up to 57 generations, an emergent path-dependent semantic selection mechanism that was not part of the original design. This mechanism is functionally equivalent to Spike-Timing-Dependent Plasticity (STDP) in biological neural circuits, and—unlike most reports of emergence in contemporary AI—its magnitude is predictable a priori from the closed-form expression $0.997^{56}\!\approx\!0.846$, which sits within the variance of the measured value $\mathrm{CR}_{2}\approx 0.905$ at the 56th generation. The remainder of this introduction sets out the theoretical and engineering background against which this finding becomes intelligible: why an algebro-deterministic substrate, rather than a probabilistic or analog one, was needed in the first place.
@@ 3.1 VaCoAl’s Algebraic Diffusion and Kanerva’s Random Projection
VaCoAl, on the basis of expert mathematical verification, invariably selects this $G(x)$ as a primitive polynomial. In rigorous mathematical terms, an $m$-th-degree polynomial $G(x)$ is primitive if and only if it satisfies all three of the following algebraic conditions:
@@ 3.2 Digital Implementation of Analog Diffusion: AD Functional Induction of Random Projection
As detailed in the preceding section, the action of the “mathematical mixer” == “Avalanche Effect” produced by the rigorously designed primitive polynomial $G(x)$ guarantees, through deterministic computation alone and without probabilistic random-number generation, that the “50% inversion rule” holds statistically. This ensures that, as the dimensionality of the ultra-high-dimensional space increases, the correlation (inner product) between generated vectors converges relatively to zero—a property mathematically guaranteed. We term this “quasi-orthogonality.”

### 16 Apr 2026 v3

@@ 3.1 VaCoAl’s Algebraic Diffusion and Kanerva’s Random Projection
Implementation note on the role of $F(x)$.
Equation (5) expresses $F(x)$ as the concatenation of an information region $x^{m}P(x)$ and a parity region $x^{m}P(x)\bmod G(x)$, after the manner of a systematic BCH code. In the PyVaCoAl implementation, however, only the parity region is materialised and consumed: the remainder is masked to $m$ bits and used directly as an SRAM/DRAM address, while the information region is never carried forward as bits because the underlying datum $P(x)$ is already indexed by the label (target_id) stored at that address. Equation (5) should therefore be read as a conceptual decomposition that motivates the algebraic structure of the transformation; the physical output of the diffusion stage is the $m$-bit residue alone. This engineering simplification is lossless with respect to the downstream majority-voting retrieval and is the reason a single primitive-polynomial LFSR per block suffices to realise the full mapping.
@@
VaCoAl selects the feedback polynomial $G_{b}(x)$ of each block-$b$ LFSR from the class of long-period polynomials of degree $64$ over $\mathrm{GF}(2)$, obtained by sampling a $64$-bit integer uniformly at random and fixing its degree-0 and degree-$64$ coefficients to $1$. Rigorously primitive polynomials of degree 6464—those satisfying irreducibility, generation of the multiplicative group of $\mathrm{GF}(2^{64})$, and maximal period $2^{64}-1$ simultaneously—are the optimal but not strictly necessary choice; empirically, the overwhelming majority of randomly sampled degree-$64$ polynomials of the above form possess periods long enough, and avalanche statistics close enough to the $50\%$ ideal, to deliver the quasi-orthogonality on which the architecture relies. The three conditions below are therefore stated as the asymptotic ideal toward which the design aspires; the PyVaCoAl implementation does not enforce them explicitly, and deviation from the ideal manifests as a small, empirically bounded degradation in the per-block error distribution that is absorbed by the majority-voting circuit of Section 3.4.
The primitive-polynomial ideal.
A degree-$m$ polynomial $G(x)\in\mathrm{GF}(2)[x]$ is called primitive if and only if the following three conditions hold simultaneously.
@@ 3.2 Digital Implementation of Analog Diffusion: AD Functional Induction of Random Projection
As detailed in the preceding section, the action of the “mathematical mixer”—the Avalanche Effect produced by the long-period degree-$64$ feedback polynomial $G_{b}(x)$—guarantees, through deterministic computation alone and without probabilistic random-number generation, that the $50\%$ inversion rule” holds statistically to within a tolerance that diminishes with $N$ and is absorbed by majority voting. This ensures that, as the dimensionality of the ultra-high-dimensional space increases, the correlation (inner product) between generated vectors converges relatively to zero—a property mathematically guaranteed. We term this “quasi-orthogonality.”
@@ 3.3 Divide and Conquer: Induction of Distributed Storage
Granularity of diffusion: global versus per-block.
For expositional clarity, Sections 3.1–3.2 have presented the Galois-field diffusion as a single global transformation $F(x)$ applied to the entire input vector, with block partitioning introduced as a subsequent step. In the actual PyVaCoAl implementation the order is reversed: the input hyper-vector is first partitioned into $N$ segments of length $q=L/N$, and each block $b$ then applies its own independent LFSR diffusion — with its own feedback polynomial $G_{b}(x)$ and its own initial seed — to the byte-packed segment assigned to it, producing an $m$-bit address. This per-block diffusion is not merely a parallelisation convenience; it is structurally necessary for the noise-filtering argument developed in Section 3.4.1. Under a hypothetical global diffusion, a single-bit perturbation of the input would, by virtue of the avalanche effect, propagate to all $N$ block addresses, corrupting every vote and destroying the majority-based recovery. Under the per-block diffusion actually implemented, such a perturbation affects only the one or two blocks whose segments contain the flipped bit; the remaining NO(1)N-O(1) blocks continue to output the correct entry address, and the avalanche effect within the affected blocks diffuses their erroneous votes uniformly across the address space, producing the “flat-field” noise distribution exploited by the Chernoff-bound argument in Appendix A. The two descriptions — the global $F(x)$ of Sections 3.1–3.2 and the per-block LFSR array of the present section — are algebraically compatible (both are deterministic linear maps over $\mathrm{GF}(2)$), but only the latter supports the locality of noise on which majority-voting noise tolerance depends.
@@ 3.3.3 Collision Tolerance (Don’t Care)
3.3.4 Concrete Structure of the Multi-Stage Rescue Circuit
When PyVaCoAl operates in Rescue mode (rescue-sample-rate $r>0$), the exogenous collision-avoidance circuit that complements the intrinsic Don’t Care mechanism is realised as a four-stage pipeline, here described at a level of detail sufficient to reproduce the behaviour.
  1. 1.
    Sample accumulation at write time. For every learned pair $(hv,\mathit{label})$, the triple $(\mathbf{a},\mathbf{s},\mathit{tid})$ is appended to a temporary buffer, where $\mathbf{a}\in\mathbb{Z}_{2^{m}}^{N}$ is the vector of per-block addresses produced by the diffusion stage, $\mathbf{s}\in\{0,1\}^{N\times\lceil q/8\rceil}$ is the byte-packed per-block segment of the input, and $\mathit{tid}$ is the integer label identifier assigned to the concept.
  2. 2.
    Per-block address-sorted finalisation. After learning completes, the buffer is re-indexed by block. For each block $b\in\{1,\dots,N\}$, the triples $\{(a_{b}^{(k)},s_{b}^{(k)},\mathit{tid}^{(k)})\}_{k}$ drawn from all samples are sorted in ascending order of $a_{b}$. The resulting per-block sorted arrays constitute a compact, DRAM-resident auxiliary table $\mathcal{R}_{b}$ of size $O(K)$ per block, where $K$ is the number of learned samples. No hash-table pointer-chasing is required; $\mathcal{R}_{b}$ is a flat, cache-coherent array.
  3. 3.
    Binary search at query time. When the main memory read returns $v_{b}<0$ for block $b$ — whether $v_{b}=-1$ (a legacy collision marker) or $v_{b}<-1$ (a bucket reference) — the query’s address $a_{b}^{q}$ is used as the search key against the sorted addresses of $\mathcal{R}_{b}$, via two searchsorted calls (left and right), yielding the half-open interval $\mathit{lo},\mathit{hi})$ of candidate entries colliding at $a_{b}^{q}$. The cost is $O(\log K)$ per block.
  4. 4.
    Segment Exact Match. Within the candidate interval, the query’s packed segment $s_{b}^{q}$ is compared byte-wise against each candidate’s stored $s_{b}^{(k)}$. The first exact match substitutes its $\mathit{tid}^{(k)}$ for the negative $v_{b}$, restoring the block to the voting pool. If no candidate matches, the block’s vote is discarded (a genuine Don’t Care). This step is critical: it preserves the invariant “one block == at most one vote” even on bucket-referenced cells that host multiple colliding $\mathit{tid}$s, and thereby safeguards the discriminative power of the majority vote.
This four-stage pipeline is the concrete realisation of the “multi-stage collision-avoidance circuit” invoked abstractly in Sections 3.1–3.2. In Rescue mode with full sampling ($r=1$), stages 1–4 deterministically recover every collided block for which a truly matching segment exists, achieving FS-relative Exact Match and pinning the block-voting rate CR1 at $1.0$ throughout — at the cost, documented in Sections 3.5 and 5.6, of annihilating the CR2-based path-discrimination capability. In Don’t Care mode ($r=0$) the pipeline is suppressed in its entirety, and the residual per-block collisions contribute the minute analog variance in CR1 whose multiplicative accumulation into CR2 drives the emergent semantic selection analysed in Section 5.6.